OpenStudy (dtan5457):

Divide 1/3(a^3b^4) by 2/3(a^2b)

OpenStudy (dtan5457):

\[\frac{ 1 }{ 3 } a^3b^4 \div \frac{ 2 }{ 3 } a^2b\]

OpenStudy (bibby):

rewrite as a multiplication of the reciprocal of the right and then cancel accordingly

OpenStudy (bibby):

\(\huge \frac{a^m}{a^n}=a^{m-n}\)

OpenStudy (bibby):

also worth noting \(\huge \frac{2}{3}a^2b=\frac{2a^2b}{3}\)

OpenStudy (triciaal):

group like terms its easier to "see" when you divide by a fraction invert the divisor and multiply rules of exponents already posted above; when you divide numbers or variables with the same base subtract the exponents

OpenStudy (dtan5457):

so basically reciprocal the 2/3 to 3/2 and then just divide normally?

OpenStudy (triciaal):

multiply

OpenStudy (bibby):

\(\large \frac{a}{b} \div \frac{c}{d}= \frac{a}{b}*\frac{d}{c}\)

OpenStudy (dtan5457):

ok, but the terms a and b, divide them right?

OpenStudy (bibby):

in this case a is (a^3b^4) and b is 3

OpenStudy (dtan5457):

1/2 ab^3

OpenStudy (bibby):

my point is \(\frac{1}{2} ab^3=\frac{ab^3}{2} \)

OpenStudy (triciaal):

then for the variables a^3/a^2 = a^(3-2) = a^1 b^4/b= b^(4-1)= b^3 so altogether 1/2ab^3 good job!

OpenStudy (bibby):

oh I'm stupid

OpenStudy (dtan5457):

i got it now

OpenStudy (dtan5457):

thanks guys

OpenStudy (triciaal):

since you are learning exponents and using same bases you might want to not work with fractions like elementary but for this problem use base of 3 1/3 = 3^-1 divide by 2*3^-1 what you have is 1/2* 3^-1-(-1) so it become 1/2 * 3^(-1 + 1) anything raised to the power 0 = 1 so 1/2 *3^0 = 1/2 *1 = 1/2